Allen bought 20 stamps at the post office in 37¢ and 20¢ denominations. If the total cost of the stamps was $7.06, how many 37¢ stamps did Allen buy?
Answer by checkley71(8403)

X+Y=20 X BEING THE NUMBER OF $.37 & Y BEING THE $.20 STAMPS.
X=20-Y
.37X+.20Y=7.06
.37(20-Y)+.20Y=7.06
7.40-.37Y+.20Y=7.06
-.17Y=7.06-7.40
-.17Y=-.34
Y=-34/-17
Y=2 NUMBER OF $.20 STAMPS.
X+2=20
X=20-2
X=18 NUMBER OF $.37 STAMPS.
PROOF
.37*18+.20*2=7.06
6.66+.40=7.06
7.06=7.06